# Chirality-Driven Magnetization Emerges from Relativistic Four-Current Dynamics

### v1

@@ Abstract
Chirality-induced spin selectivity (CISS) is a striking quantum phenomenon in which electron transport through chiral molecules leads to spin polarization—even in the absence of magnetic fields or magnetic components. Although observed in systems such as DNA, helicenes, proteins, and polymers, the fundamental physical origin of CISS remains unresolved. Here, we introduce a time-dependent relativistic four-current framework, in which charge and current densities evolve according to the time-dependent variational principle. Real-time relativistic four-current simulations enable direct analysis of helical currents and induced magnetization dynamics. Applied to helicenes—axially chiral molecules lacking stereocenters—our simulations reveal curvature-induced helical electron currents that generate spontaneous magnetic fields aligned along the molecular axis. These fields are handedness-dependent and reach magnitudes of $10^{-1}\sim 10^{-2}$ Tesla per single helicene strand. Our results suggest that CISS may arise from intrinsic, relativistic curvature-induced helical currents and the associated magnetic fields within chiral molecules. This four-current mechanism offers a self-contained explanation for spin selectivity, independent of interfacial effects or strong spin–orbit coupling. Furthermore, our results lead to several testable hypotheses that can be explored in experiments.
@@
The current density is the most important quantum mechanical property that underlies the magnetic response and dynamics of the electrons, including NMR spectroscopy, magnetic susceptibilities, spin transport, and magnetically induced current pathways in molecules and materials.95 Within the Schrödinger framework, current can be introduced as a velocity density.96, 97 However, the nonrelativistic velocity density fails to capture spin-dependent currents or those arising from spin-couplings (e.g., spin–orbit, spin–spin), such as those described by the Dirac–Coulomb–Breit operator.98, 99, 100 The full treatment of the four-current within the four-component relativistic framework was developed by Saue and co-workers for studies of magnetically induced current density,101 and was later extended to the London orbital framework.102
@@
Figure 3: Highest singly occupied molecular orbital of anionic [24]helicene: (top) field-free ground state and (bottom) under a 0.01 au electric field aligned along the $-z$-direction.
@@
For the electron transport simulations, an extra electron is added to the helicene molecule, forming an anionic molecular system. The spatial distribution of this additional electron in the ground state is shown in Fig. 3. In the absence of an external electric field, the electron is delocalized across the entire molecule, with the highest density concentrated near its center. The relativistic RT-X2C-TDDFT electron dynamics are initiated from an electric-field-polarized state (see Theory and Computational Details for more information), with the initial molecular orbital depicted in Fig. 3. The quantum dynamics are driven by an external field $E_{z}=-0.01$ au in the $-z$ direction.
@@
To examine the current vector in more detail, Fig. 5 displays the current direction in the plane perpendicular to the chiral axis (the $z$-direction). As shown, (L)-[24]helicene produces a counterclockwise current, whereas (R)-[24]helicene generates a clockwise current when viewed from above, looking down along the $-z$-axis. We would like to emphasize that in both simulations, the direction of charge transfer is identical, specifically, from $+z$ to $-z$. This suggests that the observed difference in current circulation arises solely from the molecular handedness.
@@
To investigate the length dependence of the helical-current-induced magnetic field, a series of helicenes with varying heights were simulated using four-current dynamics. The maximum magnetic field generated for each helicene is shown in Fig. 8. The field strength ranges from $10^{-1}\sim 10^{-2}$ Tesla. As evident from the figure, the magnetic field increases with helicene height, reaching a peak at [16]helicene. This trend suggests the presence of a critical helical length beyond which the helical-current-induced magnetic field saturates. This behavior can be attributed to the finite delocalization length of $\pi$-conjugated electrons, which governs low-energy electron transport in such systems.
@@ Conclusion
The computed helical-current-induced magnetic field for a single helicene lies in the range of $10^{-1}$ to $10^{-2}$ Tesla. Its magnitude increases with the length of the helicene, reaching a maximum at a critical threshold length, around [16]helicene, beyond which further increases in length do not significantly enhance the field.
The helical-current-induced magnetic field likely drives the CISS effect by breaking spin symmetry and favoring one spin orientation, enabling spin filtering.
@@
  • The helical-current-induced magnetic field can, in turn, enter the expression for the helical current itself as a vector potential in the second term of Eq. 5. Consequently, a self-consistent treatment of the field–current interaction may further amplify the observed magnetic field.
@@
In this work, we use the exact-two-component (X2C) transformation to yield an electron-only relativistic Hamiltonian.103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 75, 116, 117, 118, 119, 120, 121, 122 Specifically, we use the one-electron X2C transformation, which allows for one-step approach to construct a unitary transformation matrix to “fold” the small component coefficients into the large component. The one-electron X2C approach makes use of the effective one-electron spin–orbit Hamiltonian and avoids the four-component self-consistent-field procedure. In this work, we use the new Dirac–Coulomb–Breit-parameterized effective one-electron spin–orbit Hamiltonian in the X2C approach.123
@@
Commonly used density functional approximations that only depend on the spin-collinear densities ($\mathbf{P}^{\alpha\alpha}$ and $\mathbf{P}^{\beta\beta}$) must be modified to be compatible with two-component Hamiltonians.124, 125, 126, 127, 128, 129, 107, 130, 131, 132, 108, 133, 134, 135, 136, 137, 138, 139, 140, 141, 142, 117 The non-collinear DFT framework used in this work redefines the functionals (in the similar spirit of spin-polarized DFT) to depend on a set of auxiliary generalized variables which take the full magnetization vector density into account,134, 137, 138, 117 As this approach has been developed and calibrated, we refer readers to Ref. 117 for mathematical derivations and implementations and Ref. 143 for a recent review of non-collinear DFT methods.
@@
The Gordon decomposition of the two-component current density can be derived in the following manner. The current density is given by144
@@
The relativistic magnetic field arising from a four-current density can be obtained from Jefimenko’s equation of the magnetic field, 145, 146
@@
We constructed a periodic single helicene chain in an orthogonal lattice with lattice constants $a=20$  and $b=20$  to eliminate interchain interactions along the $x$- and $y$-directions. The helical axis is along the $z$-direction. The lattice constant along $z$ was optimized using the Vienna Ab initio Simulation Package (VASP).147 The electron–ion interaction was treated using the projector augmented-wave (PAW) method148 with a plane-wave energy cutoff of 525 eV. For the exchange–correlation functional, the Perdew–Burke–Ernzerhof (PBE) form of the generalized gradient approximation (GGA)149 was used, along with Grimme’s DFT-D3 dispersion correction.150 A $1\times 1\times 10$ Gamma-centered Monkhorst–Pack $k$- point mesh was employed. The optimized lattice constants were $a=20$, $b=20$, $c=3.515$  for both left- and right-handed helicenes.
The helicene molecule was cleaved from the periodic model and further geometry-optimized using the Gaussian 16 package with the B3LYP exchange–correlation functional and Def2-SV(P) basis set. Grimme’s DFT-D3 dispersion correction with Becke–Johnson damping (DFT-D3BJ) was also included.151 The optimized helicene exhibited no imaginary frequencies, confirming a stable minimum.
To investigate the electronic dynamics, we employed the relativistic RT-X2C-TDDFT approach using a development version of the open-source Chronus Quantum package.152, 153 The initial state ($t=0$) was prepared via a self-consistent field (SCF) calculation under a static electric field of 0.01 au along the $-z$-direction. To initialize the electronic dynamics, an electric field of equal magnitude but opposite direction was applied, and the system was propagated in real time using the RT-X2C-TDDFT method. The time evolution was carried out for 3 fs with a time step of 0.05 au At each step, the time-dependent two-component density was Gordon-decomposed into four-current. All real time calculations used the 6-311G(d,p) basis set.

### v2

@@ Abstract
Chirality-induced spin selectivity (CISS) is a striking quantum phenomenon in which electron transport through chiral molecules leads to spin polarization—even in the absence of external magnetic fields or magnetic components. Although observed in systems such as DNA, helicenes, proteins, and polymers, the fundamental physical origin of CISS remains unresolved. Here, we introduce a time-dependent relativistic four-current framework, in which charge and current densities evolve according to the time-dependent variational principle. Real-time relativistic four-current simulations enable direct analysis of helical currents and induced magnetization dynamics. Applied to helicenes—axially chiral molecules lacking stereocenters—our simulations reveal curvature-induced helical electron currents that generate spontaneous magnetic fields aligned along the molecular axis. These fields are handedness-dependent and reach magnitudes of $10^{-1}$ Tesla per single helicene strand. Our results suggest that CISS may arise from intrinsic, relativistic curvature-induced helical currents and the associated magnetic fields within chiral molecules. This four-current mechanism offers a self-contained explanation for the driving force underlying spin selectivity, independent of interfacial effects or unphysically enhanced spin–orbit coupling. Furthermore, our results provide a new perspective that offers a unifying framework with the potential to reconcile many existing hypotheses and theoretical models, while also suggesting several testable predictions that can be examined experimentally.
@@
The current density is the most important quantum mechanical property that underlies the magnetic response and dynamics of the electrons, including NMR spectroscopy, magnetic susceptibilities, spin transport, and magnetically induced current pathways in molecules and materials.95, 96, 97, 98, 99 Within the Schrödinger framework, current can be introduced as a velocity density.100, 101 However, the nonrelativistic velocity density fails to capture spin-dependent currents or those arising from spin-couplings (e.g., spin–orbit, spin–spin), such as those described by the Dirac–Coulomb–Breit operator.102, 103, 104 The full treatment of the four-current within the four-component relativistic framework was developed by Saue and co-workers for studies of magnetically induced current density,105 and was later extended to the London orbital framework.106
@@
Figure 3: Highest singly occupied molecular orbital of anionic [24]helicene: (top) field-free ground state and (bottom) under a 0.01 au electric field aligned along the z-z-direction.
@@
Specifically, we aim to simulate the experimental condition in which a helical system is subjected to an electric field bias, resulting in electron transport through the helicene.27, 29, 30, 31, 32, 33, 34 For the electron transport simulations, an extra electron is added to the helicene molecule, forming an anionic molecular system. The introduction of a spin-1/2 system with an extra electron is specifically designed to probe the spin-current contribution, which can be rigorously described within the relativistic four-current formalism.
The spatial distribution of this additional electron in the ground state is shown in Fig. 3. In the absence of an external electric field, the electron is delocalized across the entire molecule, with the highest density concentrated near its center. The relativistic RT-X2C-TDDFT electron dynamics are initiated from an electric-field-polarized state (see Theory and Computational Details for more information), with the initial molecular orbital depicted in Fig. 3. The quantum dynamics are driven by an external field $E_{z}=-0.01$ au in the z-z direction. The total simulation time is 1 fs, and a complex absorbing boundary condition was applied.107 During the simulation, the electrons did not reach the boundaries; therefore, no reflection occurred, and we observed no significant difference between the scenarios with and without the complex absorbing boundary condition. This setup represents a non-equilibrium transport condition.108, 109, 107
@@
To examine the current vector in more detail, Fig. 5 displays the current direction in the plane perpendicular to the chiral axis (the $z$-direction). As shown, (L)-[24]helicene produces a counterclockwise current, whereas (R)-[24]helicene generates a clockwise current when viewed from above, looking down along the z-z-axis. We would like to emphasize that in both simulations, the direction of charge transfer is identical, specifically, from +z+z to z-z. This suggests that the observed difference in current circulation arises solely from the molecular handedness.
@@
To investigate the length dependence of the helical-current-induced magnetic field, a series of helicenes with varying heights were simulated using four-current dynamics. The maximum magnetic field generated for each helicene is shown in Fig. 7. The field strength is $\sim 10^{-1}$ Tesla. As shown in the figure, the magnetic field increases with helicene height and reaches a plateau beyond [16]helicene. This trend suggests the presence of a critical helical length beyond which the helical-current-induced magnetic field saturates. This behavior can be attributed to the finite delocalization length of $\pi$-conjugated electrons, which governs low-energy electron transport in such systems.
@@
From the Boltzmann population analysis of the Kramers’ pair (spin-down and spin-up) at 5 K under a 0.46 T magnetic field generated by the helical current in [24]helicene, the spin-down/spin-up ratio is approximately 0.9, indicating that the magnetic field favors one spin manifold by about 10%. However, this small magnitude suggests that the helical-current-induced magnetic field, although able to break time-reversal symmetry, is by itself insufficient to account for the full extent of the CISS effect. Additional mechanisms must therefore exist to amplify the magnetic field induced by the helical current.
@@
From the perspective of a foundational theoretical framework, several important conclusions can be drawn. The velocity current (first term) can be fully accounted for within non-relativistic quantum dynamics, indicating that the observed spontaneous generation of the magnetic field has a non-relativistic origin. Although the spin-dependent paramagnetic current (third term) requires a fully variational relativistic treatment, its contribution in the helicene system is negligible, indicating that a self-consistent treatment of spin and spin–orbit interactions is not necessary in this case. However, in systems with strong spin and spin–orbit couplings, the spin-dependent paramagnetic current can become significant, demanding the use of relativistic four-current dynamics. A fully self-consistent treatment of the diamagnetic term (second term) requires a quantum field description, which could be approximated using a semiclassical approach in future work.
@@ Conclusion
The computed helical-current-induced magnetic field for a single helicene lies in the range of $10^{-1}$ Tesla. Its magnitude increases with the length of the helicene, reaching a maximum at a critical threshold length, around [16]helicene, beyond which further increases in length do not significantly enhance the field.
This work examines the origin of spontaneous magnetization arising from helicity-induced electron currents. As shown in Eq. 6, if the four-current does not have a preferential axis or is not associated with a principal axis of electron transport, the isotropic averaging of the induced field vanishes, resulting in no net magnetization. In contrast, when electron transport occurs along a well-defined direction in a helical structure, the resulting preferential alignment of the four-current breaks isotropic cancellation and gives rise to a non-vanishing spontaneous magnetization. This highlights the fundamental connection between helicity, directional electron motion, and the breaking of isotropic symmetry, which together underpin the emergence of magnetization in chiral systems. The helical-current-induced magnetic field is likely the origin of the CISS effect by breaking spin symmetry and favoring one spin orientation, enabling spin filtering.
While the system studied here is specifically tied to helicity-induced electron currents, the four-current quantum dynamics framework is completely general. It enables the analysis of a broad range of driving forces in diverse physical and chemical contexts, including systems with intrinsic chiral centers, strong spin–orbit interactions, and other symmetry-breaking mechanisms.
@@
The physical insights gained from this work provide a new perspective on the CISS effect. While the helical-current-induced magnetization accounts for spin Zeeman splitting, it must be amplified to explain the full extent of spin selectivity. This consideration introduces the concepts of a ‘driving force’ and ‘amplification’ in CISS, offering a unifying framework with the potential to reconcile many of the existing hypotheses and theoretical models. The ‘driving force’ is the helical current, which may arise from electron transport through a helical structure and/or from chiral phonons, the latter also correlating with temperature effects. Potential ‘amplification’ mechanisms include strong spin–orbit coupling, which enhances spin-dependent paramagnetic contributions; molecular aggregation, which increases the overall current; and quantum-field self-consistent interactions, which strengthen the diamagnetic response.
@@
  • The helical-current-induced magnetic field can, in turn, enter the expression for the helical current itself as a vector potential in the second term of Eq. 5. Consequently, a self-consistent treatment of the field–current interaction may further amplify the observed magnetic field. Ideally, these effects would be rigorously treated within a quantum field framework, providing motivation for the development of four-current quantum dynamics coupled with quantum-field-enabled time-dependent electronic structure theory.
@@ Methods
In this work, we use the exact-two-component (X2C) transformation to yield an electron-only relativistic Hamiltonian.110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 75, 123, 124, 125, 126, 127, 128, 129 Specifically, we use the one-electron X2C transformation, which allows for one-step approach to construct a unitary transformation matrix to “fold” the small component coefficients into the large component. The one-electron X2C approach makes use of the effective one-electron spin–orbit Hamiltonian and avoids the four-component self-consistent-field procedure. In this work, we use the new Dirac–Coulomb–Breit-parameterized effective one-electron spin–orbit Hamiltonian in the X2C approach.130 The relativistic Hamiltonian does not include explicit quantum field operators; therefore, the four-current formalism employed in this work should be regarded as an effective quantum electrodynamics treatment. It is important to note that the perturbative Pauli Hamiltonian is not bounded from below and, therefore, is unsuitable for describing relativistic four-current dynamics.
@@
Commonly used density functional approximations that only depend on the spin-collinear densities ($\mathbf{P}^{\alpha\alpha}$ and $\mathbf{P}^{\beta\beta}$) must be modified to be compatible with two-component Hamiltonians.131, 132, 133, 134, 135, 136, 114, 137, 138, 139, 115, 140, 141, 142, 143, 144, 145, 146, 147, 148, 149, 124 The non-collinear DFT framework used in this work redefines the functionals (in the similar spirit of spin-polarized DFT) to depend on a set of auxiliary generalized variables which take the full magnetization vector density into account,141, 144, 145, 124 As this approach has been developed and calibrated, we refer readers to Ref. 124 for mathematical derivations and implementations and Ref. 150 for a recent review of non-collinear DFT methods.
@@
The Gordon decomposition of the two-component current density can be derived in the following manner. The current density is given by151
@@
The relativistic magnetic field arising from a four-current density can be obtained from Jefimenko’s equation of the magnetic field, 152, 153
@@
We constructed a periodic single helicene chain in an orthogonal lattice with lattice constants $a=20$ Å  and $b=20$ Å  to eliminate interchain interactions along the $x$- and $y$-directions. The helical axis is along the $z$-direction. The lattice constant along $z$ was optimized using the Vienna Ab initio Simulation Package (VASP).154 The electron–ion interaction was treated using the projector augmented-wave (PAW) method155 with a plane-wave energy cutoff of 525 eV. For the exchange–correlation functional, the Perdew–Burke–Ernzerhof (PBE) form of the generalized gradient approximation (GGA)156 was used, along with Grimme’s DFT-D3 dispersion correction.157 A $1\times 1\times 10$ Gamma-centered Monkhorst–Pack $k$- point mesh was employed. The optimized lattice constants were $a=20$ Å, $b=20$ Å, $c=3.515$ Å  for both left- and right-handed helicenes.
The helicene molecule was cleaved from the periodic model and further geometry-optimized using the Gaussian 16 package with the B3LYP exchange–correlation functional and Def2-SV(P) basis set. Grimme’s DFT-D3 dispersion correction with Becke–Johnson damping (DFT-D3BJ) was also included.158 The optimized helicene exhibited no imaginary frequencies, confirming a stable minimum.
To investigate the electronic dynamics, we employed the relativistic RT-X2C-TDDFT approach using a development version of the open-source Chronus Quantum package.159, 160 The initial state ($t=0$) was prepared via a self-consistent field (SCF) calculation under a static electric field of 0.01 au along the z-z-direction. To initialize the electronic dynamics, an electric field of equal magnitude but opposite direction was applied, and the system was propagated in real time using the RT-X2C-TDDFT method. The time evolution was carried out for 3 fs with a time step of 0.05 au At each step, the time-dependent two-component density was Gordon-decomposed into four-current. All real time calculations used the 6-311G(d,p) basis set.
1 Code Availability
The real-time X2C-TDDFT framework is implemented in the public version of the Chronus Quantum software package.159 All calculations were performed using executables compiled with GCC version 13.2.
2 Data Availability
[…]
3 Author Contributions
S. Upadhyay, X. Zheng, and X. Li conceived the project and developed the computer code. S. Upadhyay, X. Zheng, T. Wang, and A. Shayit carried out the calculations and analyzed the data. J. Liu, D. Sun, and X. Li acquired research funding. S. Upadhyay, X. Zheng, and X. Li wrote the manuscript with input from all authors. All authors discussed the results and approved the final version of the manuscript.